OpenStudy (anonymous):

(-6m^7 + 20m^6) ÷ 2m^4 A. -3m7 + 10m6 B. -6m7 + 20m2 C. -3m3 + 10m2 D. -3m3 + 20m6

OpenStudy (anonymous):

I think it is A

OpenStudy (anonymous):

if not a then C

OpenStudy (unklerhaukus):

\[(-6m^7 + 20m^6) ÷ 2m^4\]\[=\frac{(-6m^7 + 20m^6)} {2m^4}\]\[=\frac{2m^4(-3m^3+10m^2)}{2m^4}\]